What's the wavelength of a 400-nm wave with 3*10^8 speed on light?
Answer #1

wavelength is 400nm

400nm = 400 *10^-9

frequency = speed / wavelength = 3* 10^8 / 400 *10^-9 = 7.5 * 10^14 Hz